Determine the measure of an angle whose measure is 2/3 of its complement
Answer by jorel1380(3719)

1/(2/3)n=3n/2
n+3/2n=90
5/2n=90
5n=180
n=36
3/2n=54
36 degrees is 2/3 of a 54 degree angle; which is its complement..